# SEAL: Searching Expandable Architectures for Incremental Learning

### v1

@@ 4.6 The novel figure of merits of SEAL
$\mathcal{F}(\bar{x},\vec{d},w)=(1-\mathrm{AA_{K}}(\bar{x},\vec{d}))\cdot\texttt{size}(\bar{x},\vec{d})^{w},\ \ w=0.07$
@@
For the second objective, we use a rescaling-invariant flatness metric, a similar measurement used in [11]. In this definition, the flatness is measured as the change in performance when perturbing the weights under multiplicative noise. Given $\sigma$ the noise intensity, the final flatness is evaluated using a neighborhood of $N$ (i.e., sampling N perturbations):
$\Delta\mathrm{ACC}_{K}=\|\mathrm{ACC}_{\mathcal{D}_{K}}(W_{K})-\mathrm{ACC}_{\mathcal{D}_{K}}(W_{K}+\sigma\cdot z_{i}\odot W_{K})\|$
$\mathcal{H}(\bar{x},\vec{d},\sigma)=\dfrac{1}{N}\sum_{i=1}^{N}\dfrac{\mathrm{ACC}_{\mathcal{D}_{K}}(W_{K})-\Delta\mathrm{ACC}_{K}(W_{K},\sigma)}{\mathrm{ACC}_{\mathcal{D}_{K}}(W_{K})}$
@@ 5.1 Datasets
We evaluate our methods on three benchmark datasets: CIFAR-10 [19], CIFAR-100 [19], and ImageNet16-120[5]. CIFAR-10 and CIFAR-100 consist of $32\times 32$ color images, representing 10 and 100 classes, respectively. Each dataset contains 60000 images split into 50000 training images and 10000 testing images. ImageNet16-120 is a downsized version of the ImageNet dataset [6] tailored for NAS evaluation, comprising $16\times 16$ color images across 120 object classes.
@@ 5.2 Analysis of the results
For all the experiments, the MobileNetV3-based [13] supernet is used and the hyperparameters of the NN architecture set $\Omega_{\bar{x}}$ consist of resolution, depth, width, and kernel size. The value of the threshold $\tau$ for expansion is 0.2 for CIFAR-10, and 0.13 fo the other datasets. Although the most relevant metric is the $AA$ computed after the last sample set training, all the metrics are tracked across the training. Tables 1 and 2 report the average accuracy, average forgetting, and forward transfer values. The average incremental accuracy and the backward transfer are not presented as their values directly relate to the average accuracy and the average forgetting, respectively. Finally, to evaluate the models post-search, we aim at simulating a real-case scenario. As in incremental learning we are interested in the change in performance for the training samples. Instead of evaluating the model in a completely new dataset we augment the original training set with more samples, coming from data splits not available during the NAS. In this way, we are able to accurately measure how the model and expansion policy will perform in the future. These experiments were performed under a regime of K=5 sequential splits of disjoint samples, where the model was trained with 3 epochs per task. We emphasize that the model is expanded only if triggered by the condition presented in 4.3, checked on every new task instead of growing at any task to retain the model complexity.
@@
Regarding the baselines, we followed the benchmarking done in [32] and selected models that suited our constraints and our data-incremental setting. Based on these criteria, we selected EWC [18], LWF [21], and SI [37]— being straightforward yet state-of-the-art methods across a range of continual learning settings. To strengthen the baselines, we included the naive and the joint-training benchmark cases, common baselines in IL. The naive method is a model trained in an incremental way without any additional constraint, it shows the performance degradation of a model just by being trained in a continual setting. On the other hand, the joint-training refers to a model trained on the whole dataset at once, as if it were a static dataset. SEAL has been executed with 5 different seeds on CIFAR-10 and CIFAR-100 and once on Imagenet16-120. All the benchmarks have been trained with 5 different seeds. In Table 1, we show our main result, i.e., that our SEAL models are competitive with the other baselines in every metric on each dataset. Notably, the naive approach sometimes achieves high accuracy—especially on CIFAR-10—when tasks are similar and do not require strong regularization. However, while SEAL outperforms naive in overall performance, it also shows increased forgetting and lower forward transfer on CIFAR-10, potentially indicating overfitting on the latest tasks. Finally, joint-training does not always yield the highest accuracy, likely due to task interference in heterogeneous datasets.
@@
Naive 93.18 (±$(\pm 0.05)$ 1.41 (±$(\pm 0.50)$ 0.68 (±$(\pm 0.27)$
Joint 94.85 (±$(\pm 0.15)$
EWC 93.87 (±$(\pm 0.61)$ 1.13 (±$(\pm 0.30)$ 1.33 (±$(\pm 0.45)$
SI 94.31 (±$(\pm 0.37)$ 0.71 (±$(\pm 0.43)$ 1.44 (±$(\pm 0.02)$
LWF 92.49 (±$(\pm 2.16)$ 2.18 (±$(\pm 2.27)$ 1.22 (±$(\pm 0.84)$
Ours 95.35 (±$(\pm 0.18)$ 1.76 (±$(\pm 0.41)$ 0.18 (±$(\pm 0.19)$
@@
Naive 81.32 (±$(\pm 0.85)$ 7.65 (±$(\pm 0.66)$0.45 (±$(\pm 1.21)$
Joint 84.71 (±$(\pm 0.65)$
EWC 82.38 (±$(\pm 0.37)$ 6.94 (±$(\pm 0.40)$ 0.58 (±$(\pm 0.60)$
SI 82.55 (±$(\pm 0.39)$ 5.89 (±$(\pm 0.48)$0.53 (±$(\pm 0.75)$
LWF 83.21 (±$(\pm 0.25)$ 5.50 (±$(\pm 0.14)$ 2.65 (±$(\pm 0.27)$
Ours 83.94 (±$(\pm 0.33)$ 4.81 (±$(\pm 0.37)$ 1.59 (±$(\pm 0.25)$
@@
Naive 51.71 (±$(\pm 0.15)$ 5.54 (±$(\pm 0.35)$ 1.00 (±$(\pm 0.37)$
Joint 51.84 (±$(\pm 1.30)$
EWC 51.75 (±$(\pm 0.03)$ 5.50 (±$(\pm 0.65)$ 2.00 (±$(\pm 1.14)$
SI 51.39 (±$(\pm 0.12)$ 5.73 (±$(\pm 0.44)$ 0.83 (±$(\pm 0.41)$
LWF 53.02 (±$(\pm 0.10)$ 4.10 (±$(\pm 0.15)$ 3.27 (±$(\pm 0.24)$
Ours 52.30 (±$(\pm 0)$ 3.35 (±$(\pm 0)$ 6.91 (±$(\pm 0)$
An initial experiment compares performance between single-objective optimization using the mono-objective genetic algorithm Differential Evolution [30], and multi-objective optimization incorporating the flatness objective. Table 2 shows metrics for four models, each row aggregating the searched model from five NAS runs: the searched single-objective model (accuracy) and the searched multi-objective models by accuracy, trade-off, and flatness.
CIFAR-10@5
Selection Params Flatness ACC Forgetting FWT
SingleObj 3.98M (±$(\pm 0.63)$ 73.61 (±$(\pm 1.82)$ 94.82 (±$(\pm 0.57)$ 2.09 (±$(\pm 1.00)$ 0.55 (±$(\pm 0.74)$
MultiObj-Acc 3.92M (±$(\pm 0.50)$ 75.05 (±$(\pm 1.14)$ 95.35 (±$(\pm 0.18)$ 1.76 (±$(\pm 0.41)$ 0.18 (±$(\pm 0.19)$
MultiObj-Flat 3.79M (±$(\pm 0.40)$ 82.09 (±$(\pm 2.13)$ 94.53 (±$(\pm 0.51)$ 1.37 (±$(\pm 0.54)$ 0.75 (±$(\pm 0.34)$
MultiObj-Trade 3.65M (±$(\pm 0.15)$ 78.19 (±$(\pm 3.33)$ 94.89 (±$(\pm 0.68)$ 1.89 (±$(\pm 0.56)$ 0.25 (±$(\pm 0.25)$
CIFAR-100@5
Selection Params Flatness ACC Forgetting FWT
SingleObj 4.37M (±$(\pm 0.06)$ 52.57 (±$(\pm 1.84)$ 83.01 (±$(\pm 1.20)$ 5.72 (±$(\pm 1.28)$ 0.83 (±$(\pm 1.02)$
MultiObj-Acc* 4.09M (±$(\pm 0.21)$ 57.88 (±$(\pm 3.48)$ 83.94 (±$(\pm 0.33)$ 4.81 (±$(\pm 0.37)$ 1.59 (±$(\pm 0.25)$
MultiObj-Flat 3.85M (±$(\pm 0.02)$ 58.49 (±$(\pm 2.87)$ 83.61 (±$(\pm 0.01)$ 5.14 (±$(\pm 0.04)$ 1.12 (±$(\pm 0.22)$
MultiObj-Trade* 4.09M (±$(\pm 0.21)$ 57.88 (±$(\pm 3.48)$ 83.94 (±$(\pm 0.33)$ 4.81 (±$(\pm 0.37)$ 1.59 (±$(\pm 0.25)$
ImageNet-16@5
Selection Params Flatness ACC Forgetting FWT
SingleObj* 4.24M (±$(\pm 0.51)$ 45.42 (±$(\pm 1.82)$ 52.31 (±$(\pm 0.01)$ 3.35 (±$(\pm 0.28)$ 6.91 (±$(\pm 0.17)$
MultiObj-Acc* 4.24M (±$(\pm 0.51)$ 45.42 (±$(\pm 1.82)$ 52.31 (±$(\pm 0.01)$ 3.35 (±$(\pm 0.28)$ 6.91 (±$(\pm 0.17)$
MultiObj-Flat 3.55M (±$(\pm 0.06)$ 56.16(±$(\pm 1.89)$ 50.40 (±$(\pm 0.20)$ 2.29(±$(\pm 0.22)$ 4.21 (±$(\pm 0.34)$
MultiObj-Trade 4.06M (±$(\pm 0.36)$ 49.98 (±$(\pm 2.59)$ 51.51 (±$(\pm 0.29)$ 2.24 (±$(\pm 0.26)$ 4.33(±$(\pm 0.25)$
@@ 6 Conclusions
The aim of this paper was to propose SEAL, the first NAS framework able to optimize jointly the architecture and the expansion policy for a DNN. Our results show that the learned policies effectively reduce forgetting, improve the model performance, and reduce the memory complexity compared to previous methods. SEAL has been designed for a data-incremental setting but could be extended to many other relevant IL settings. Another promising direction lies in enhancing the role of NAS within the model growing policy. In this work, certain decisions were treated as fixed hyperparameters based on the assumption that architectures would behave similarly across the splits of our data-incremental setup. However, this assumption may not hold in domain-incremental or class-incremental scenarios. A possible improvement would be to incorporate a capacity-related control signal (e.g., a threshold-bit) directly into the NAS process, allowing the search to determine when and how the model should expand dynamically.

### v2

@@ 4.6 The novel figure of merits of SEAL
$\mathcal{F}(\bar{x},\vec{d},w)=(1-\mathrm{AA_{K}}(\bar{x},\vec{d}))\cdot\texttt{size}(\bar{x},\vec{d})^{w},\ \ w=0.07$ (12)
@@
For the second objective, we use a rescaling-invariant flatness metric, a similar measurement used in (Gambella, Pittorino, and Roveri 2024). In this definition, the flatness is measured as the change in performance when perturbing the weights under multiplicative noise. Given $\sigma$ the noise intensity, the final flatness is evaluated using a neighborhood of $N$ (i.e., sampling N perturbations):
$\Delta\mathrm{ACC}_{K}$ $=\left\|\mathrm{ACC}_{\mathcal{D}_{K}}(W_{K})-\mathrm{ACC}_{\mathcal{D}_{K}}(W_{K}+\sigma\cdot z_{i}\odot W_{K})\right\|$ (13)
$\mathcal{H}(\bar{x},\vec{d},\sigma)$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{\mathrm{ACC}_{\mathcal{D}_{K}}(W_{K})-\Delta\mathrm{ACC}_{K}(W_{K},\sigma)}{\mathrm{ACC}_{\mathcal{D}_{K}}(W_{K})}$ (14)
@@ 5.1 Datasets
We evaluate our methods on three benchmark datasets: CIFAR-10 (Krizhevsky 2009), CIFAR-100 (Krizhevsky 2009), and ImageNet16-120(Chrabaszcz, Loshchilov, and Hutter 2017). CIFAR-10 and CIFAR-100 consist of $\times 32$ 32 color images, representing 10 and 100 classes, respectively. Each dataset contains 60000 images split into 50000 training images and 10000 testing images. ImageNet16-120 is a downsized version of the ImageNet dataset (Deng et al. 2009) tailored for NAS evaluation, comprising $\times 16$ 16 color images across 120 object classes.
@@ 5.2 Analysis of the results
For all the experiments, the MobileNetV3-based (Howard 2017) supernet is used and the hyperparameters of the NN architecture set $\Omega_{\bar{x}}$ consist of resolution, depth, width, and kernel size. The value of the threshold $\tau$ for expansion is 0.2 for CIFAR-10, and 0.13 fo the other datasets. Although the most relevant metric is the $AA$ computed after the last sample set training, all the metrics are tracked across the training. Tables 2 and 1 report the average accuracy, average forgetting, and forward transfer values. The average incremental accuracy and the backward transfer are not presented as their values directly relate to the average accuracy and the average forgetting, respectively. Finally, to evaluate the models post-search, we aim at simulating a real-case scenario. As in incremental learning we are interested in the change in performance for the training samples. Instead of evaluating the model in a completely new dataset we augment the original training set with more samples, coming from data splits not available during the NAS. In this way, we are able to accurately measure how the model and expansion policy will perform in the future. These experiments were performed under a regime of K=5 sequential splits of disjoint samples, where the model was trained with 3 epochs per task. We emphasize that the model is expanded only if triggered by the condition presented in 4.3, checked on every new task instead of growing at any task to retain the model complexity.
@@
Regarding the baselines, we followed the benchmarking done in (van de Ven, Tuytelaars, and Tolias 2022) and selected models that suited our constraints and our data-incremental setting. Based on these criteria, we selected EWC (Kirkpatrick et al. 2016), LWF (Li and Hoiem 2017), and SI (Zenke, Poole, and Ganguli 2017)— being straightforward yet state-of-the-art methods across a range of continual learning settings. To strengthen the baselines, we included the naive and the joint-training benchmark cases, common baselines in IL. The naive method is a model trained in an incremental way without any additional constraint, it shows the performance degradation of a model just by being trained in a continual setting. On the other hand, the joint-training refers to a model trained on the whole dataset at once, as if it were a static dataset. SEAL has been executed with 5 different seeds on all the datasets. All the benchmarks have been trained with 5 different seeds. In Table 2, we show our main result, i.e., that our SEAL models are competitive with the other baselines in every metric on each dataset. Notably, the naive approach sometimes achieves high accuracy—especially on CIFAR-10—when tasks are similar and do not require strong regularization. However, while SEAL outperforms naive in overall performance, it also shows increased forgetting and lower forward transfer on CIFAR-10, potentially indicating overfitting on the latest tasks. Finally, joint-training does not always yield the highest accuracy, likely due to task interference in heterogeneous datasets.
An initial experiment compares performance between single-objective optimization using the mono-objective genetic algorithm Differential Evolution (Price, Storn, and Lampinen 2005), and multi-objective optimization incorporating the flatness objective. Table 1 shows metrics for four models, each row aggregating the searched model from five NAS runs: the searched single-objective model (accuracy) and the searched multi-objective models by accuracy, trade-off, and flatness.
CIFAR-10@5
Selection Params Flatness ACC Forgetting FWT
SingleObj 3.98M (±0.63)($\pm$ 0.63)( ± 0.63 ) 73.61 (±1.82)($\pm$ 1.82)( ± 1.82 ) 94.82 (±0.57)($\pm$ 0.57)( ± 0.57 ) 2.09 (±1.00)($\pm$ 1.00)( ± 1.00 ) 0.55 (±0.74)($\pm$ 0.74)( ± 0.74 )
MultiObj-Acc 3.92M (±0.50)($\pm$ 0.50)( ± 0.50 ) 75.05 (±1.14)($\pm$ 1.14)( ± 1.14 ) 95.35 (±0.18)($\pm$ 0.18)( ± 0.18 ) 1.76 (±0.41)($\pm$ 0.41)( ± 0.41 ) 0.18 (±0.19)($\pm$ 0.19)( ± 0.19 )
MultiObj-Flat 3.79M (±0.40)($\pm$ 0.40)( ± 0.40 ) 82.09 (±2.13)($\pm$ 2.13)( ± 2.13 ) 94.53 (±0.51)($\pm$ 0.51)( ± 0.51 ) 1.37 (±0.54)($\pm$ 0.54)( ± 0.54 ) 0.75 (±0.34)($\pm$ 0.34)( ± 0.34 )
MultiObj-Trade 3.65M (±0.15)($\pm$ 0.15)( ± 0.15 ) 78.19 (±3.33)($\pm$ 3.33)( ± 3.33 ) 94.89 (±0.68)($\pm$ 0.68)( ± 0.68 ) 1.89 (±0.56)($\pm$ 0.56)( ± 0.56 ) 0.25 (±0.25)($\pm$ 0.25)( ± 0.25 )
CIFAR-100@5
Selection Params Flatness ACC Forgetting FWT
SingleObj 4.37M (±0.06)($\pm$ 0.06)( ± 0.06 ) 52.57 (±1.84)($\pm$ 1.84)( ± 1.84 ) 83.01 (±1.20)($\pm$ 1.20)( ± 1.20 ) 5.72 (±1.28)($\pm$ 1.28)( ± 1.28 ) 0.83 (±1.02)($\pm$ 1.02)( ± 1.02 )
MultiObj-Acc* 4.09M (±0.21)($\pm$ 0.21)( ± 0.21 ) 57.88 (±3.48)($\pm$ 3.48)( ± 3.48 ) 83.94 (±0.33)($\pm$ 0.33)( ± 0.33 ) 4.81 (±0.37)($\pm$ 0.37)( ± 0.37 ) 1.59 (±0.25)($\pm$ 0.25)( ± 0.25 )
MultiObj-Flat 3.85M (±0.02)($\pm$ 0.02)( ± 0.02 ) 58.49 (±2.87)($\pm$ 2.87)( ± 2.87 ) 83.61 (±0.01)($\pm$ 0.01)( ± 0.01 ) 5.14 (±0.04)($\pm$ 0.04)( ± 0.04 ) 1.12 (±0.22)($\pm$ 0.22)( ± 0.22 )
MultiObj-Trade* 4.09M (±0.21)($\pm$ 0.21)( ± 0.21 ) 57.88 (±3.48)($\pm$ 3.48)( ± 3.48 ) 83.94 (±0.33)($\pm$ 0.33)( ± 0.33 ) 4.81 (±0.37)($\pm$ 0.37)( ± 0.37 ) 1.59 (±0.25)($\pm$ 0.25)( ± 0.25 )
ImageNet-16@5
Selection Params Flatness ACC Forgetting FWT
SingleObj* 4.24M (±0.51)($\pm$ 0.51)( ± 0.51 ) 45.42 (±1.82)($\pm$ 1.82)( ± 1.82 ) 52.31 (±0.01)($\pm$ 0.01)( ± 0.01 ) 3.35 (±0.28)($\pm$ 0.28)( ± 0.28 ) 6.91 (±0.17)($\pm$ 0.17)( ± 0.17 )
MultiObj-Acc* 4.24M (±0.51)($\pm$ 0.51)( ± 0.51 ) 45.42 (±1.82)($\pm$ 1.82)( ± 1.82 ) 52.31 (±0.01)($\pm$ 0.01)( ± 0.01 ) 3.35 (±0.28)($\pm$ 0.28)( ± 0.28 ) 6.91 (±0.17)($\pm$ 0.17)( ± 0.17 )
MultiObj-Flat 3.55M (±0.06)($\pm$ 0.06)( ± 0.06 ) 56.16(±1.89)($\pm$ 1.89)( ± 1.89 ) 50.40 (±0.20)($\pm$ 0.20)( ± 0.20 ) 2.29(±0.22)($\pm$ 0.22)( ± 0.22 ) 4.21 (±0.34)($\pm$ 0.34)( ± 0.34 )
MultiObj-Trade 4.06M (±0.36)($\pm$ 0.36)( ± 0.36 ) 49.98 (±2.59)($\pm$ 2.59)( ± 2.59 ) 51.51 (±0.29)($\pm$ 0.29)( ± 0.29 ) 2.24 (±0.26)($\pm$ 0.26)( ± 0.26 ) 4.33(±0.25)($\pm$ 0.25)( ± 0.25 )
@@
Naive 93.18 (±0.05)($\pm$ 0.05)( ± 0.05 ) 1.41 (±0.50)($\pm$ 0.50)( ± 0.50 ) 0.68 (±0.27)($\pm$ 0.27)( ± 0.27 )
Joint 94.85 (±0.15)($\pm$ 0.15)( ± 0.15 ) - -
EWC 93.87 (±0.61)($\pm$ 0.61)( ± 0.61 ) 1.13 (±0.30)($\pm$ 0.30)( ± 0.30 ) 1.33 (±0.45)($\pm$ 0.45)( ± 0.45 )
SI 94.31 (±0.37)($\pm$ 0.37)( ± 0.37 ) 0.71 (±0.43)($\pm$ 0.43)( ± 0.43 ) 1.44 (±0.02)($\pm$ 0.02)( ± 0.02 )
LWF 92.49 (±2.16)($\pm$ 2.16)( ± 2.16 ) 2.18 (±2.27)($\pm$ 2.27)( ± 2.27 ) 1.22 (±0.84)($\pm$ 0.84)( ± 0.84 )
Ours 95.35 (±0.18)($\pm$ 0.18)( ± 0.18 ) 1.76 (±0.41)($\pm$ 0.41)( ± 0.41 ) 0.18 (±0.19)($\pm$ 0.19)( ± 0.19 )
@@
Naive 81.32 (±0.85)($\pm$ 0.85)( ± 0.85 ) 7.65 (±0.66)($\pm$ 0.66)( ± 0.66 ) -0.45 (±1.21)($\pm$ 1.21)( ± 1.21 )
Joint 84.71 (±0.65)($\pm$ 0.65)( ± 0.65 ) - -
EWC 82.38 (±0.37)($\pm$ 0.37)( ± 0.37 ) 6.94 (±0.40)($\pm$ 0.40)( ± 0.40 ) 0.58 (±0.60)($\pm$ 0.60)( ± 0.60 )
SI 82.55 (±0.39)($\pm$ 0.39)( ± 0.39 ) 5.89 (±0.48)($\pm$ 0.48)( ± 0.48 ) -0.53 (±0.75)($\pm$ 0.75)( ± 0.75 )
LWF 83.21 (±0.25)($\pm$ 0.25)( ± 0.25 ) 5.50 (±0.14)($\pm$ 0.14)( ± 0.14 ) 2.65 (±0.27)($\pm$ 0.27)( ± 0.27 )
Ours 83.94 (±0.33)($\pm$ 0.33)( ± 0.33 ) 4.81 (±0.37)($\pm$ 0.37)( ± 0.37 ) 1.59 (±0.25)($\pm$ 0.25)( ± 0.25 )
@@
Naive 51.71 (±0.15)($\pm$ 0.15)( ± 0.15 ) 5.54 (±0.35)($\pm$ 0.35)( ± 0.35 ) 1.00 (±0.37)($\pm$ 0.37)( ± 0.37 )
Joint 51.84 (±1.30)($\pm$ 1.30)( ± 1.30 ) - -
EWC 51.75 (±0.03)($\pm$ 0.03)( ± 0.03 ) 5.50 (±0.65)($\pm$ 0.65)( ± 0.65 ) 2.00 (±1.14)($\pm$ 1.14)( ± 1.14 )
SI 51.39 (±0.12)($\pm$ 0.12)( ± 0.12 ) 5.73 (±0.44)($\pm$ 0.44)( ± 0.44 ) 0.83 (±0.41)($\pm$ 0.41)( ± 0.41 )
LWF 53.02 (±0.10)($\pm$ 0.10)( ± 0.10 ) 4.10 (±0.15)($\pm$ 0.15)( ± 0.15 ) 3.27 (±0.24)($\pm$ 0.24)( ± 0.24 )
Ours 52.45 (±0.13)($\pm$ 0.13)( ± 0.13 ) 3.23 (±0.24)($\pm$ 0.24)( ± 0.24 ) 6.98 (±0.18)($\pm$ 0.18)( ± 0.18 )
@@ 6 Conclusions
This paper introduced SEAL, the first NAS framework to jointly optimize both architecture and expansion policy for DNNs. Our results demonstrate that SEAL reduces forgetting, improves performance, and lowers memory complexity compared to prior methods. While designed for data-incremental learning, SEAL can be extended to other IL settings. A promising direction for future work is to strengthen NAS’s role in dynamic model growth. In this study, some decisions were fixed under the assumption that architectures generalize across data splits—an assumption that may not hold in domain- or class-incremental settings. Incorporating a capacity-aware control signal (e.g., a threshold-bit) into the NAS process could enable more adaptive expansion decisions.